OpenStudy (anonymous):

evaluate 8cos((5pi)/6)

OpenStudy (anonymous):

For this questions I need to find the position of the mass at time t. x(t) = 8 cos t. The value of t is given as 5pi/6

OpenStudy (anonymous):

8cos150 = -8cos30 = -4sqrt3

OpenStudy (anonymous):

cos(5*pi / 6) is equal to -0.866, so: x(5*pi/6) = 8 * (-0.866) = - 6.92

OpenStudy (anonymous):

Thank you both! Do you have a good resource to review the unit circle and the values? I found the decimal answer, but I believe the online homework wants me to include the radical answers.

OpenStudy (anonymous):

Although when I put it in the calculator, I got positive 6.9282

OpenStudy (souvik):

you can do it using this... \[\cos5\pi/6=\cos(\pi-\pi/6)=-\cos \pi/6\]

OpenStudy (anonymous):

I need to remember to figure the quadrant into the answer! Second quadrant means my answer would be negative. Thank you!

OpenStudy (dan815):